OpenStudy (anonymous):

Find the coordinates of the center of the following circle. (x - 3)² + (y - 7)² = 49

OpenStudy (anonymous):

(3,7)

OpenStudy (anonymous):

The general equation of a circle is : \[(x-a)^2+(y-b)^2=r^2\] Where (a,b) us the coordinates of the center and r is the radius

OpenStudy (anonymous):

thank you !!(:

OpenStudy (anonymous):

so the center points are (a,b)

OpenStudy (anonymous):

why wouldnt it be (-3,-7)

OpenStudy (anonymous):

@Shianne95 Look at what I have written above : \[(x−a)^2+(y−b)^2=r^2\] So the center is (a,b) not (-a,-b)

OpenStudy (anonymous):

okay thanks